OpenStudy (anonymous):

Solve 3^(x – 4) = 7^(x + 9)

OpenStudy (anonymous):

start with \[(x-4)\ln(3)=(x+9)\ln(7)\] then a bunch of algebra to solve for \(x\)

OpenStudy (anonymous):

@satellite73 ok? would the answer be x ≈ –25.86

OpenStudy (anonymous):

i have no idea, you have to do some algebra first

OpenStudy (anonymous):

@satellite73 I'm stuck on the algebra, can you help me work it out?

OpenStudy (anonymous):

remember that \(\ln(3)\) and \(\ln(7)\) are just numbers

OpenStudy (anonymous):

distribute first \[\ln(3)x-4\ln(3)=\ln(7)x+9\ln(7)\]

OpenStudy (anonymous):

then put everything with an \(x\) on one side, everything else on the other \[\ln(3)x-\ln(7)x=9\ln(7)+4\ln(3)\]

OpenStudy (anonymous):

factor the \(x\) on the left \[\left(\ln(3)-\ln(7)\right)x=9\ln(7)+4\ln(3)\]

OpenStudy (anonymous):

then divide \[x=\frac{9\ln(7)+4\ln(3)}{\ln(3)-\ln(7)}\]

OpenStudy (anonymous):

i have no idea what this number is, now is when you use a calculator

OpenStudy (sirm3d):

\(x\approx-25.855890260549890173158545490824\)